solve the system of equations 2r+2s=50 and 2r-s=17
Answer by checkley77(12844)

2r+2s=50
2r-s=17 subtract
------------------------
3s=33
s=33/3
s=11 ans.
2r+2*11=50
2r+22=50
2r=50-22
2r=28
r=28/2
r=14 ans.
Proof:
2*14-11=17
28-11=17
17=17